OpenStudy (anonymous):

An urn contains 4 red balls and 6 blue balls. A second urn contains 16 red balls and an unknown amount of blue balls. If a single ball is drawn from each urn, the probability that both balls are the same color is 0.44. How many blue balls are in the second urn?

OpenStudy (anonymous):

harder than i first thought, lets see if we can do it put the number of blue balls as \(x\) then the probability the first ball and the second ball is blue is \[.6\times \frac{x}{16+x}\] and the probability that the first and second ball chosen is red is \[.4\times \frac{16-x}{16+x}\] the total is \(0.44\) so we can set \[\frac{.6x}{16+x}+\frac{.4(16-x)}{16+x}=0.44\] and try to solve for \(x\)

OpenStudy (anonymous):

oh no that was wrong sorry

OpenStudy (anonymous):

the probability that both chosen are red is \[.4\times \frac{16}{16+x}\]

OpenStudy (anonymous):

so set \[\frac{.6x}{16+x}+\frac{.4\times 16}{16+x}=0.44\]

OpenStudy (anonymous):

\[\frac{.6x+6.4}{16+x}=0.44\] \[.6x+6.4=0.44(16+x)\] etc